OpenStudy (anonymous):

how many numbers of 7 digits can be formed from the digits 1,2,3,4,5,6,7 if there are not more than 2 digits between 1 and 2?

OpenStudy (ash2326):

There are 3 cases a)1 and 2 are together b) 1 and 2 have one digit in between c) 1 and 2 have 2 digits in between Case a 1 and 2 are together so we have 6 digits now with 1 and 2 together=6!* 2! Case b 1 and 2 have one digit in between 1 _ 2 this blank can be filled in 5C1 ways 5!*5C1*2! Case c 1 and 2 have two digits in between 1 _ _2 these blanks can be filled in 5C2 ways with the digits interchanging 5C2*2! total numbers =4!*5C2*2!*2! So total ways= 6!* 2!+ 5!*5C1*2!+ 4!*5C2*2!*2! =720*2+120*5*2+24*10*2*2 =1440+1200+960 =2640+960 =3600 ways

OpenStudy (anonymous):

hey ash are you still there?

OpenStudy (anonymous):

could you please explain case b) and case c) i don't understand how you got 5!*5C1*2! and 5C2*2!

OpenStudy (anonymous):

AAHHHHH I GET IT, I UNDERSTAND NOW hehehehe THANK YOU for the very detailed explanation!!!!!!